HUBBLE SPACE TELESCOPE DAILY REPORT #1064 PERIOD COVERED: 0700 EST Daily Status Report as of 031/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 5196 (GHRS SMOV Side 1 High Voltage Turn On) The High Resolution Spectrograph (Side One) was was turned on using its high voltage for early calibrations tests. These are the first steps in the activation of the GHRS (Side One) since the repair kit was installed. during the servicing mission. Calibration dark images were also taken. The COSTAR mirrors were seen to go to the correct positions. The observations were received and displayed. They all were reported to look normal. 1.2 Completed WF/PC-2 5549 (GHRS Side 2 Dark Noise Monitor - Cycle 4) The WF/PC-2 was used to make internal tests for routine monitoring of detector dark noise. There was no carousel motion. A total of sixteen observations were received and displayed. They all were reported to be nominal. 1.3 Completed WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) The WF/PC-2 was used to make a set of visual internal flat images. Four images were received and displayed. Two of the observations were reported to be underexposed, and the other two were nominal. 1.4 Completed Two Sets of WF/PC-2 4721 (WF/PC-2 Decontamination and TEC Turn On) The WF/PC-2 was used to make ultraviolet photometric observations of the star standard A+81D262, plus K-Spots, and dark calibration images. All four sets of CCDs were used. A total of 11 images were received and displayed in the OSS. No anomalies were reported. 1.5 Completed Two Sets of WF/PC-2 4763 (WF/PC-2 QE and Throughput Measurement) The WF/PC-2 was used to make ultraviolet exposures of the UV flux-standard star BD+75D325 to be used for alignment studies. The OSS monitored the observations, and assessed the data quality. No anomalies were noted. 1.6 Completed Two Sets of S/C 5254 (GHRS SMOV Side One Activation) The High Resolution spectrograph (Side One) low voltage turned on as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 4749 (WF/PC-2 Relative Photometric Calibration) The WF/PC-2 was used to make photometric calibrations images of standard star A+81D266 which is located in NGC-5136 (aka, Omega Centauri) to obtain various medium and broad band filter images. The OSS monitored the observations and assessed the data quality. An STR track change caused the loss of data for one observation, however the target star was in the unaffected part of the image. 1.8 Completed COSTAR/U 5605 (COSTAR/FOS Fine Alignment X-Tilt/Y-Tilt (Repeat of 4747 With A New Target)) Both the Faint Object Spectrograph and the COSTAR were used to make observations of standard star WD0501+527 to perform alignment and focus operations on the FOS M1 mirror. The guide star reacquisition obtained fine lock and the observations continued without apparent anomaly. The observations were received and displayed. They all were reported to look nominal. The COSTAR/FOS scheduled mirror move was canceled, because it was not needed. 1.9 Completed FOC/96 5152 (Shape and Expansion of SN1987A Ejecta) The Faint Object Camera (f/96) was used to make observations of the ejecta from supernova SN1987A and monitor its evolution. Two different filters were used. The observations were executed in fine lock. All of the data was received in the OSS, and was reported to look fine, though the supernova remnant structure is rather faint especially in shorter exposures. 1.10 Completed WF/PC-2 5457 (A Snapshot Survey for Blue Straggles in Magellanic Cluster Cores) The WF/PC-2 was used to make observations of the populous cluster NGC-2193 which is located in the Large Magellanic cloud looking for blue straggler stars. These populous clusters are the Magellanic Cloud's equivalent of a galactic globular cluster. Two short exposures were obtained, each using a different filter. Both observations were received and displayed in the OSS. 1.11 Completed S/C 4435 (Spacecraft Disturbance Test) This engineering proposal was used to gather data to determine if the spacecraft is shaking during day/night and night/day transitions with the new solar arrays being installed. Data was obtained at two different orientations. The test was completed without apparent problems. 1.12 Complete COSTAR/FOS 5604 (COSTAR/FOS Fine Alignment and Focus) The scheduled Y-Tilt a and X-Tilt adjustment for the COSTAR FOS mirror was completed as scheduled with no problems. The observations were received and displayed in real time. A subsequent reacquisition was not successful (See Operations Note 2.1). The Take Data Flag remained down, and no observations were made. Subsequent reacquisitions were successful, However, there were indications that the target was no well centered. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 6 Successful : 5 FGS Reacquisition failure at 031/0334Z effected the FOS 5604 observations #3 through #8. An ARU/PRT was executed, and the subsequent reacquisition was successful. 2.2 FHST Updates: Scheduled :0 Successful :0 2.3 Significant HSTARs: None 2.4 Operations Notes: SMOV activities continue to be executed on or ahead of the milestone schedule. During last week, WF/PC-2 executed decontamination operations, flat fields calibration, QE and throughput measurements and point spread function characterization. FOC completed all of its SMOV activities and, along with WF/PC-2, is now performing "normal" science. Initial COSTAR/GHRS coarse alignment adjustments were made, and further GHRS side 1 operations were performed. Another target was utilized in the COSTAR/FOS fine alignment process. All HST subsystem activities were performed nominally. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WF/PC-2 5240 (SMOV WF/PC-2 Flat Fields Calibration) 3.2 Complete FOS/BL/COSTAR 5604 (COSTAR/FOS Fine Alignment - Focus Repeat of 4745 With New Target) 3.3 Complete WF/PC-2 5475 (A Snapshot Survey for Blue Stragglers in Magellanic Cluster Cores) 3.4 Complete WF/PC-2 4709 (WF/PC-2 Internal Monitor) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-